OpenStudy (perl):

A day care program has an average daily expense of $75.00. The standard deviation is $15.00. The owner takes a sample of 64 bills. What is the probability the mean of his sample will be between $70.00 and $80.00? @brittneeharley

OpenStudy (perl):

The first thing you want to do is find the z score, right?

OpenStudy (perl):

Keep in mind we are looking at the sampling distribution of sample means . So we have to use standard error , standard deviation / sqrt( n ) $$ \Large z = \frac{ 70 - 75 }{ \frac{15 }{ \sqrt{64}}}$$

OpenStudy (perl):

so thats the first z score

OpenStudy (perl):

$$ \Large{ z = \frac{ 70 - 75 }{ \frac{15 }{ \sqrt{64}}}= -2.667 \\~\\ z = \frac{ 80 - 75 }{ \frac{15 }{ \sqrt{64}}}= 2.667 } $$

OpenStudy (perl):

so you need to find the area